OpenStudy (anonymous):

Which method is correct when finding the product of (x − 3)2? Justify your answer in complete sentences. Method 1 (x − 3^)2 (x − 3)(x − 3) x2 − 6x + 9 method 2 (x − 3)^2 x^2 − 3^2 x^2 − 9

OpenStudy (anonymous):

It is the first one because the power of 2 is attached to the whole (x-3).

OpenStudy (anonymous):

The first one...

OpenStudy (anonymous):

Method 1 is the correct method

OpenStudy (anonymous):

\[(x-3)^2 = (x-3)(x-3) = x^2 + 9 - 6x\]